# An Attack on $p$-adic Lattice Public-key Cryptosystems and Signature Schemes

### v1

@@ 3.2 LVP in General Local Fields
Let $K$ be an extension field of of $\mathbb{Q}_{p}$ degree $n=ef$, where $e$ is the ramification index and $f$ is the residue degree. Let $\pi$ be a uniformizer of $K$, then $\left|\pi\right|_{p}=p^{-\frac{1}{e}}$. We can choose $f$ unities $s_{1},s_{2},\dots,s_{f}\in K$ such that
@@
Let $V$ be a left vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be a basis of $V$ over $\mathbb{Q}_{p}$ such that $\left|\alpha_{1}\right|=\left|\alpha_{2}\right|=\cdots=\left|\alpha_{n}\right|$, which is denoted by $\lambda_{1}$. Then $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ is an orthogonal basis of $V$ over $\mathbb{Q}_{p}$ if and only if
@@
Suppose that $1\leq k<l$ and $\beta_{1},\beta_{2},\dots,\beta_{k}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k})$. According to Lemma 2.4, we can check whether $\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1})$ by verifying that if
@@
Since $\left|\beta_{s}^{\prime}\right|_{p}\geq\left|\alpha_{m}\right|_{p}>\left|p\alpha_{1}\right|_{p}=\left|p\beta_{1}\right|_{p}$, we have $\lambda_{2}=\left|\beta_{s}^{\prime}\right|_{p}$ and a second longest vector is $v=\beta_{s}^{\prime}$. ∎
We summarize the algorithm in above theorem as follows.
Algorithm (finding $\lambda_{2}$).
Input: a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$.
Output: the value $\lambda_{2}$ and a lattice vector $v\in\mathcal{L}$ such that $\left|v\right|_{p}=\lambda_{2}$.
  1. 1.
    rearrange $\beta_{1},\beta_{2},\dots,\beta_{m}$ such that $\left|\beta_{1}\right|_{p}=\max{\left\{\left|\beta_{1}\right|_{p},\left|\beta_{2}\right|_{p},\dots,\left|\beta_{m}\right|_{p}\right\}}$,
  2. 2.
    $\beta_{1}^{\prime}\leftarrow\beta_{1}$, $L\leftarrow[\beta_{1}]$ (a list, indices begin with 1),
  3. 3.
    for $i=2$ to $m$ do:
  4. 4.
    $t\leftarrow{\rm length}(L)$,
  5. 5.
    for $(j_{1},j_{2},\dots,j_{t})\in\{0,1,\dots,p-1\}^{t}$ do:
  6. 6.
    if $\left|\beta_{i}-\sum_{k=1}^{t}{j_{k}L[k]}\right|_{p}<\left|\beta_{1}\right|_{p}$ then:
  7. 7.
    $\beta_{i}^{\prime}\leftarrow\beta_{i}-\sum_{k=1}^{t}{j_{k}L[k]}$, back to step 3 with the next $i$,
  8. 8.
    $\beta_{i}^{\prime}\leftarrow\beta_{i}$, append $\beta_{i}$ to $L$,
  9. 9.
    let $s$ be the subscript such that $\left|\beta_{s}^{\prime}\right|_{p}=\max{\left\{\left|\beta_{i}^{\prime}\right|_{p}:\beta_{i}\not\in L\right\}},$
Return $\lambda_{2}\leftarrow\left|\beta_{s}^{\prime}\right|_{p}$, $v\leftarrow\beta_{s}^{\prime}$.
This algorithm takes at most $(m-1)p^{f}$ many $p$-adic absolute value computations of elements of $K$. Comparing with the LVP algorithm given in [1, 2], which takes $O(p^{m})$ many $p$-adic absolute value computations, our algorithm is much faster when $f$ is much less than $m$.
@@ 4.1 The Attack
Our attack is exaclty based on recovering a uniformizer and finding an orthogonal basis. To break the schemes, we first recover a uniformizer from the public key and then use it to generate an orthogonal basis, as presented in the following lemma and theorem.
@@ 4.2 A More Effective Way
where $a_{i}\in\mathbb{Z}_{p}$ $(0\leq i\leq n-1)$. Let $L$ be the split field of $F$ over $\mathbb{Q}_{p}$. Then for each $\zeta_{i}$ ($1\leq i\leq n$), there is a Galois automorphism $\sigma_{i}\in{\rm Gal}(L/\mathbb{Q}_{p})$ which maps $\zeta$ to $\zeta_{i}$, i.e.,
@@ 4.3 A Toy Example
is the minimal polynomial of $\zeta$ and $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ are expressed as polynomials in $\zeta$ of degree less than $n$ with coefficients in $\mathbb{Z}_{2}$. He runs the algorithm “finding $\lambda_{2}$” with the input $F(x)$. Since $\left|\zeta^{i}\right|_{2}=1$ for $i=0,1,\dots,19$ and
@@
[…]
Next, Oscar runs the algorithm “CVP with orthogonal bases” in [5] with the orthogonal basis $1,\gamma,\gamma^{2},\dots,\gamma^{19}$, the lattice $\mathcal{L}(\beta_{1},\beta_{2},\beta_{3},\beta_{4})$ and the target vector $C$. The algorithm outputs a lattice vector $v=-\beta_{1}+\beta_{2}+\beta_{4}$ which is closest to $C$. Therefore, the plaintext is
@@ 5 Conclusion
Our attack extremely reduces the security strength of the $p$-adic lattice public-key encryption cryptosystems and signature schemes. It succeeds beacuse the extension field used in the signature scheme and the public-key encryption cryptosystem is totally ramified. In order to aviod this attack, we suggest to add residue degree to the extension field. As shown in Section 3.2, our attack is not efficient when the extension field has a large residue degree.

### v2

@@ 3.2 LVP in General Local Fields
Let $K$ be an extension field of of $\mathbb{Q}_{p}$ degree $n=ef$, where $e$ is the ramification index and $f$ is the residue degree. Let $\pi$ be a uniformizer of $K$, then $\left|\pi\right|_{p}=p^{-\frac{1}{e}}$. According to Proposition 6.4.5 in [7], we can choose $f$ units $s_{1},s_{2},\dots,s_{f}\in\mathcal{O}_{K}$ such that
@@
Let $V$ be a vector space over $\mathbb{Q}_{p}$ of finite dimension $n>0$, and let ||$\cdot$|| ⋅ | be a norm on $V$. Let $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ be a basis of $V$ over $\mathbb{Q}_{p}$ such that $\left|\alpha_{1}\right|=\left|\alpha_{2}\right|=\cdots=\left|\alpha_{n}\right|$, which is denoted by $\lambda_{1}$. Then $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ is an orthogonal basis of $V$ over $\mathbb{Q}_{p}$ if and only if
@@
Suppose that $1\leq k<l$ and $\beta_{1},\beta_{2},\dots,\beta_{k}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k})$. According to Lemma 3.4, $\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1})$ if and only if
$\left|\sum^{k+1}_{i=1}b_{i}\beta_{i}\right|_{p}=\lambda_{1}$
for all
$(b_{1},b_{2},\dots,b_{k+1})\in\{0,1,\dots,p-1\}^{k+1}-(0,0,\dots,0).$
Since $\beta_{1},\beta_{2},\dots,\beta_{k}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k})$, the case $b_{k+1}=0$ is already held. If $b_{k+1}\neq 0$, then, by multiplying $\left|b^{-1}_{k+1}\right|_{p}=1$ to the left side of the equation, we may assume that $b_{k+1}=1$.
Therefore, we can check whether $\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1}$ is an orthogonal basis of $\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{k},\beta_{k+1})$ by checking whether
@@
Since $\left|\beta_{s}^{\prime}\right|_{p}\geq\left|\alpha_{m}\right|_{p}>\left|p\alpha_{1}\right|_{p}=\left|p\beta_{1}\right|_{p}$, by a generalization of Lemma 3.1, we conclude that $\lambda_{2}=\left|\beta_{s}^{\prime}\right|_{p}$ and a second longest vector is $v=\beta_{s}^{\prime}$. ∎
We summarize the algorithm in above theorem as Algorithm 3.
Algorithm 3 finding $\lambda_{2}$
1:a lattice $\mathcal{L}=\mathcal{L}(\beta_{1},\beta_{2},\dots,\beta_{m})$
2:the value $\lambda_{2}$ and a lattice vector $v\in\mathcal{L}$ such that $\left|v\right|_{p}=\lambda_{2}$
3:rearrange $\beta_{1},\beta_{2},\dots,\beta_{m}$ such that $\left|\beta_{1}\right|_{p}=\max{\left\{\left|\beta_{1}\right|_{p},\left|\beta_{2}\right|_{p},\dots,\left|\beta_{m}\right|_{p}\right\}}$
4:$\beta_{1}^{\prime}\Leftarrow\beta_{1}$, $L\Leftarrow[\beta_{1}]$ (a list, indices begin with 1)
5:for $i=2$ to $m$ do
6:     $t\Leftarrow{\rm length}(L)$
7:     for $(j_{1},j_{2},\dots,j_{t})\in\{0,1,\dots,p-1\}^{t}$ do
8:         if $\left|\beta_{i}-\sum_{k=1}^{t}{j_{k}L[k]}\right|_{p}<\left|\beta_{1}\right|_{p}$ then
9:              $\beta_{i}^{\prime}\Leftarrow\beta_{i}-\sum_{k=1}^{t}{j_{k}L[k]}$
10:              back to step 3 with the next $i$
11:         end if
12:         $\beta_{i}^{\prime}\Leftarrow\beta_{i}$
13:         append $\beta_{i}$ to $L$
14:     end for
15:end for
16:let $s$ be the subscript such that $\left|\beta_{s}^{\prime}\right|_{p}=\max{\left\{\left|\beta_{i}^{\prime}\right|_{p}:\beta_{i}\not\in L\right\}}$
17:return $\lambda_{2}\Leftarrow\left|\beta_{s}^{\prime}\right|_{p}$, $v\Leftarrow\beta_{s}^{\prime}$
This algorithm takes at most $(m-1)p^{f}$ many $p$-adic absolute value computations of elements of $K$. Compared with the LVP algorithm given in [1, 2], which takes $O(p^{m})$ many $p$-adic absolute value computations, our algorithm is much faster when $f$ is much less than $m$.
@@ 4.1 The Attack
Our attack is exactly based on recovering a uniformizer and finding an orthogonal basis. To break the schemes, we first recover a uniformizer from the public key and then use it to generate an orthogonal basis, as presented in the following lemma and theorem.
@@ 4.2 A More Effective Way
where $a_{i}\in\mathbb{Z}_{p}$ $(0\leq i\leq n-1)$. Let $L$ be the splitting field of $F$ over $\mathbb{Q}_{p}$. Then for each $\zeta_{i}$ ($1\leq i\leq n$), there is a Galois automorphism $\sigma_{i}\in{\rm Gal}(L/\mathbb{Q}_{p})$ which maps $\zeta$ to $\zeta_{i}$, i.e.,
@@ 4.3 A Toy Example
is the minimal polynomial of $\zeta$ and $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ are expressed as polynomials in $\zeta$ of degree less than $n$ with coefficients in $\mathbb{Z}_{2}$. He runs Algorithm 1 with the input $F(x)$. Since $\left|\zeta^{i}\right|_{2}=1$ for $i=0,1,\dots,19$ and
@@
$C$ is also expressed as polynomials in $\zeta$ and Oscar does not know the above summands. To recover the plaintext from $C$, Oscar first expresses $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ and $C$ as polynomials in $\gamma$. The representations of $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ and $C$ are displayed in Appendix B.
Next, Oscar runs the algorithm “CVP with orthogonal basis” in [5] with the orthogonal basis $1,\gamma,\gamma^{2},\dots,\gamma^{19}$, the lattice $\mathcal{L}(\beta_{1},\beta_{2},\beta_{3},\beta_{4})$ and the target vector $C$. The algorithm outputs a lattice vector $v=-\beta_{1}+\beta_{2}+\beta_{4}$ which is closest to $C$. Therefore, the plaintext is
@@ 5 Conclusion
Our attack extremely reduces the security strength of the $p$-adic lattice public-key encryption cryptosystems and signature schemes. It succeeds because the extension field used in the signature scheme and the public-key encryption cryptosystem is totally ramified. In order to avoid this attack, we suggest to add residue degree to the extension field. As shown in Section 3.2, our attack is not efficient when the extension field has a large residue degree. According to [2], the LVP is also easy for an unramified extension, since we can choose $p\alpha_{i}$ as a second longest vector if the basis vector $\alpha_{i}$ is of length $\lambda_{1}$. Therefore, we recommend that the extension field should have large ramification index and residue degree which are almost the same.
@@
[…]
Acknowledgements
This work was supported by National Natural Science Foundation of China(No. 12271517) and National Key R&D Program of China(No. 2020YFA0712300).
Statements and Declarations
The author has no relevant financial or non-financial interests to disclose.
Appendix A Proof for the efficiency
Lemma A.1.
Write
$\zeta=a_{0}+a_{1}\theta+a_{2}\theta^{2}+\cdots+a_{n-1}\theta^{n-1}$
Proof.
$\mathbb{Z}_{p}[\zeta]=\mathbb{Z}_{p}[\zeta-a_{0}]=\mathbb{Z}_{p}[\theta].$
Lemma A.2.
For any $t=H(M\|r)$, we have $\left|t-v\right|_{p}<1$.
Proof.
Since $K=\mathbb{Q}_{p}(\theta)$, we can write
$t=a_{1}\alpha_{1}+a_{2}\alpha_{2}+\cdots+a_{n}\alpha_{n}$
Appendix B Parameters in the toy example
[…]